OpenStudy (anonymous):

Using the Product Rule, find the derivative of the following: f(x)=(x^3-3x+4)(2x-1)

OpenStudy (anonymous):

@Callisto can you help

OpenStudy (anonymous):

\[f'(x)= (3x-3)(2x-1)+(x ^{3}-3x+4)(2x)=6x ^{2}-9x+3+2x ^{4}-6x ^{2}+8x=2x ^{4}-9x+3\]

OpenStudy (anonymous):

you can finish combining the like terms.

OpenStudy (anonymous):

can you explain so i know how to do it on my own next time

OpenStudy (anonymous):

well we use the product rule and that states that f'(x) = the derivative of the 1st function times the 2nd function plus the derivative of the 2nd function times the 1st function.

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

yes.... so for the next problem i would do \[(3x ^{2})(x ^{2}-1)+(x ^{3})(2x)\]

OpenStudy (anonymous):

whats the next problem?

OpenStudy (anonymous):

that seems right if the problem is f(x)=x^3(x^2-1)

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

yup thank you mucho